OpenStudy (anonymous):

A craft store bought 27 1/4 ounces of potpourri. The shop divides the potpourri into 2 1/2 ounces bags and sells each bag for $4.25. A. How many full 2 1/2 ounces bags of potpourri can the shop make? Show work. B. If all the bags are sold, how much money will the shop bring in? Show your work. Medal, Follow, and Testimonial

OpenStudy (texaschic101):

(27 1/4) / (2 1/2) = (109/4) / (5/2) = 109/4 * 2/5 = 218/20 = 10 9/10 (thats 10 and 9/10) shop can make 10 full bags if each full bag is sold for 4.25 each, the shop will make 42.50

OpenStudy (anonymous):

Okay! Thank you so much for your help @texaschic101